OpenStudy (sh3lsh):

Stars 'n' bars Theorem Problem. 1. A bagel shop has 8 kinds of bagels. How many ways to buy a dozen bagels? 2. A bagel shop has 8 kinds of bagels. How many ways to buy a dozen bagels, with at most 4 onion and at most 2 poppy seed?

OpenStudy (anonymous):

do you know Combinations and Permutations

OpenStudy (anonymous):

12C1 +12C2 +12C3+...12C8

OpenStudy (sh3lsh):

Oh, are you sure? I know the answer for both problems, I'm just looking for the methods. The first solution should have k=4;n=8 .

OpenStudy (zarkon):

are these your answers 50388 and 35846?

OpenStudy (sh3lsh):

Let me run it through Wolfram Alpha, I just have the combinatoric form.

OpenStudy (sh3lsh):

Unfortunately, no to the first one. I have 11C4, which is 165. For the second, it was correct!

OpenStudy (zarkon):

35846 is correct but 50388 is not?

OpenStudy (sh3lsh):

Honestly, perhaps I'm concluding the wrong answer for the first question, we didn't complete the problem to its entirety. I just have k =4; n=8, which in the stars n bars theorem is just (10 choose 8). Perhaps it meant something else.

OpenStudy (sh3lsh):

We just wrote that 8 bagels were determined, so the question boiled down to x1+x2...+x8=4

OpenStudy (zarkon):

shouldn't you have \[{8-1+12\choose 8-1}\] or \[{8-1+12\choose 12}\]

OpenStudy (zarkon):

if you are using stars and bars

OpenStudy (sh3lsh):

My understanding that it's a little different with a lower and upper bound. I don't understand the concepts yet to agree nor disagree with what you wrote.

OpenStudy (zarkon):

you have 8 different bagels ||||||| the bars make 8 possible locations or bagels 1||||||| |2|||||| ||3||||| |||4|||| ||||5||| |||||6|| ||||||7| |||||||8

OpenStudy (zarkon):

let xxxxxxxxxxxx be the bagels

OpenStudy (zarkon):

you want to arrange |||||||xxxxxxxxxxxx

OpenStudy (zarkon):

the above is an example of all 12 bagels of the same kind (the 8th kind of bagel)

OpenStudy (zarkon):

the number of ways to arrange it is \[{7+12\choose 7}={8-1+12\choose 8-1}\]

OpenStudy (zarkon):

anyway...the number of ways to do problem 1 has to be larger than the number of was to do #2

OpenStudy (zarkon):

since we are putting on a restriction

OpenStudy (sh3lsh):

Ah! Okay!

OpenStudy (sh3lsh):

You're correct, I wrote my response referring a different question. I agree with what you wrote!

OpenStudy (zarkon):

so you know how to get the 35846

OpenStudy (zarkon):

?

OpenStudy (sh3lsh):

I know how to get the first answer. For the second, would we use the inclusion - exclusion principle?

OpenStudy (zarkon):

yes

OpenStudy (sh3lsh):

All Solutions - solutions with x> 4 - solutions with y> 2 + intersection where (x> 4 and y> 2)? How do I find all solutions?

OpenStudy (zarkon):

that is it we already did...\[{8-1+12\choose 8-1}=50388 \]

OpenStudy (sh3lsh):

That's obvious. I don't even know why I asked that question. I got it from here , you're the bomb. Thanks so much.

OpenStudy (zarkon):

no problem